OpenStudy (anonymous):

In a large bag of Skittles®, each of the 5 colours (red, orange, yellow, blue, green) occurs the same probability. You reach in, and select 1 candy. Find P(orange or red)

OpenStudy (anonymous):

NOTE; THIS IS AN APEX QUESTION.

OpenStudy (valpey):

P(orange or red) = P(orange) + P(red) P(red)= P(orange)= P(yellow)= P(blue)= P(green) P(red)+P(orange)+P(yellow)+P(blue)+P(green)=1

OpenStudy (valpey):

You can also solve this with a simplifying assumption. Suppose this large bag of Skittles® contains exactly five candies; one each of red, orange, yellow, blue, and green.

OpenStudy (anonymous):

So the answer is 1? I'm actually terrible at this. >_<

OpenStudy (valpey):

P(something) means "the probability or chance that something occurs". We say P(x) = 1 or 100% if it is certain to occur. If I flip a coin, P(Heads) = 0.5. P(roll a 4) of a six-sided die is 1/6 or 0.1666..

OpenStudy (valpey):

When I say "P(red)+P(orange)+P(yellow)+P(blue)+P(green)=1" I am saying that we are certain that one of the five possibilities will be true for sure.

OpenStudy (valpey):

When I say "P(red)= P(orange)= P(yellow)= P(blue)= P(green)" I am saying that the chance of picking any specific color is the same as any other specific color. (Unlike with M&Ms, say where you are more likely to pick a brown M&M).

OpenStudy (valpey):

When I say "P(orange or red) = P(orange) + P(red)" I am pointing out that these are two distinct outcomes so the chance of either one happening is equal to the sum of the chances of each one happening separately. This is because you are only picking one candy so you can't pick both a orange and a red candy at the same time. This is different from something like P(I am sick tomorrow or It will snow tomorrow) is not equal to P(I am sick tomorrow) + P(It will snow tomorrow) because it is possible that I am sick and it is snowing.

OpenStudy (valpey):

P(x) is always just a real number between 0.0 and 1.0.